OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0

OpenStudy (campbell_st):

is it 1. \[\cos^2(x) - \cos(x) = 0\] or 2. \[co(2x) - \cos(x) = 0\]

OpenStudy (anonymous):

2nd one

OpenStudy (freckles):

use a double angle identity

OpenStudy (anonymous):

ok Thank you